# The (R)Evolution of Multimodal Large Language Models: A Survey

### v1

@@ 3.2 Image Generation and Editing
While initial MLLMs demonstrated proficiency in extracting information from visual data, contemporary research has expanded its focus to include the generation of visual outputs. This advancement is realized through integrating MLLM frameworks with image generation mechanisms, predominantly embodied by the Stable Diffusion (SD) Rombach et al. (2022) models. These models feature a denoising U-Net Ronneberger et al. (2015) architecture conditioned on textual or visual embeddings, through cross-attention layers. A complete list of the analyzed models is presented in Table 3.
@@
A different approach has been studied by Li et al. (2023e) which proposes to fine-tune the LLM by adding two special tokens (i.e., <start> and <end>), and directly encode the generated text between these two tokens using the text encoder in the SD model. Similarly, in Xia et al. (2023a) the LLM is trained to output detailed language-based generation prompts which are employed for generation or editing tasks. The U-Net is fine-tuned with longer and more detailed textual captions. Furthermore, in DreamLLM Dong et al. (2023) an alignment loss is eschewed in favor of a score distillation loss while keeping the U-Net frozen. Additional research endeavors have been conducted to introduce MLLMs in the field of image editing Fu et al. (2024); Huang et al. (2023b); Tang et al. (2023).
@@ 3.3 Other Modalities and Applications
Video Understanding. Although much of the research focuses on images, some works propose MLLMs specifically designed to handle video sequences. These models process video frames independently, using CLIP-based backbones to extract frame-level features which are then combined using pooling mechanisms Li et al. (2023i); Maaz et al. (2023) or Q-Former based solutions Li et al. (2023g); Ren et al. (2023a). The connection between visual features and the language model mainly follows the same trend as image-based MLLMs, with linear projections being the most common choice. However, there are also some attempts to develop video-specific adapters Liu et al. (2023g); Ma et al. (2023a) that can capture fine-grained temporal information. In addition to encoding video frames, some works Munasinghe et al. (2023); Zhang et al. (2023b) also employ audio features to enrich the representation of input video sequences.
Any-Modality Models. Almost all models described so far treat a single modality as input to the LLM. However, a significant body of work focuses on designing effective solutions that can handle multiple modalities. This is usually achieved by aligning multimodal features through Transformer blocks such as Q-Former Chen et al. (2023c); Panagopoulou et al. (2023) and Perceiver Zhao et al. (2023d), or by utilizing ImageBind Girdhar et al. (2023) to effectively extract features that are inherently multimodal Su et al. (2023). Images, videos, and audios are the most commonly treated modalities. Additionally, some works also effectively encode 3D data Yin et al. (2023c) and IMU sensor signals Moon et al. (2023). While all these solutions can manage multimodal inputs, approaches like NExT-GPT Wu et al. (2023b) and Unified-IO 2 Lu et al. (2023a) are also capable of generating outputs of different modalities.
Domain-Specific MLLMs. In addition to dealing with generic visual inputs, some research efforts are dedicated to developing MLLMs for specific domains and applications, either training the model starting from a pre-trained LLM or fine-tuning an existing MLLM with domain-specific data. Some examples are MLLMs designed for document analysis and text-intensive visual inputs Lv et al. (2023); Ye et al. (2023a), those proposed for embodied AI and robotics Driess et al. (2023); Mu et al. (2023), and those tailored for specific domains such as medicine Li et al. (2023c) and autonomous driving Xu et al. (2023b). A complete list of domain-specific MLLMs is reported in the supplementary.
@@ 4 Conclusion and Future Directions
In this survey, we have provided a comprehensive overview of the recent evolution of MLLMs, first focusing on how to equip LLMs with multimodal capabilities and then exploring the main tasks addressed by these models. Based on the analysis presented, in the following we outline important open challenges and promising future research directions to further empower MLLMs.
Correction of Hallucinations. Several studies Liu et al. (2023b); Zhu et al. (2023a) show that MLLMs tend to exhibit high hallucination rates, especially when generating longer captions. While some solutions are emerging to mitigate this problem Liu et al. (2023b); Wang et al. (2023a); Wu et al. (2023c); Yin et al. (2023a), understanding and correcting the underlying causes of hallucinations remains an important open challenge that is worth addressing to allow the application of these models in more critical contexts (e.g., medicine) and guarantee their accuracy and trustworthiness.
Prevent Harmful and Biased Generation. Ensuring the safety and fairness of large-scale models is of fundamental interest in the community. Recent works show that models trained on web-crawled data are prone to generate inappropriate and biased content. Although recent efforts are being made to reduce this phenomenon in text-to-image generative models Schramowski et al. (2023); Friedrich et al. (2023), further exploration is needed to prevent the same behavior in MLLMs Pi et al. (2024).
@@
Acknowledgements
@@
Appendix A Additional Training Data
@@ Appendix B Evaluation
B.1 Standard Benchmarks
One of the most important skills of MLLMs is their ability to effectively answer questions based on the given input image. This ability is quantitatively evaluated across several visual question answering datasets, measuring the accuracy (Antol et al., 2015) of the answers provided by the MLLM.
@@ B.1 Standard Benchmarks
B.2 MLLM-Specific Benchmarks
@@
POPE Li et al. (2023j) is a valuable benchmark for evaluating object hallucination challenges within MLLMs. This dataset encompasses several distinct subsets, namely random, popular, and adversarial, which are generated utilizing a variety of sampling methodologies. Cumulatively, it is a binary classification query dataset that comprises 8,910 entries, facilitating comprehensive investigations into the phenomenon of object hallucination within the context of MLLMs.
@@
SEED-Bench (SEED) Li et al. (2023b) is specifically designed to evaluate LLMs and MLLMs across 12 dimensions spanning from scene understanding to OCR and action recognition. The benchmark consists of 19k multiple-choice questions written by human annotators.
@@
B.3 Visual Grounding Evaluation
@@
Visual Genome Krishna et al. (2017) connects structured image concepts to language and comprises 108,077 images along with detailed descriptions of all objects present in them, providing 5.4M region descriptions and 1.7M visual question-answer pairs. This dataset is typically used for region-level captioning and question answering.
Visual7W Zhu et al. (2016) is a visual question answering dataset that combines textual descriptions with image regions through object-level grounding. It comprises 328k question-answer pairs on 47k COCO images, together with 1.3M human-generated multiple-choice and more than 560k object groundings from 36,579 categories.
@@
B.4 Image Generation and Editing Evaluation
@@
PaLI Chen et al. (2023i) TPUv4 1,024
@@
Appendix C Computational Requirements
@@
Appendix D Further Details on Video-based and Domain-Specific MLLMs
As a complement of Sec. 3.3, we report in Table 12 and Table 13 a summary of the main characteristics of video-based and domain-specific MLLMs, respectively. For each model, we indicate the LLM used as starting point, which in some cases is initialized with the parameters of a pre-trained MLLM, the visual encoder, and the main tasks and capabilities of the MLLM. Additionally, we specify whether the LLM is kept frozen, is entirely fine-tuned, or is fine-tuned with PEFT-based strategies.
@@ Appendix D Further Details on Video-based and Domain-Specific MLLMs
Appendix E Interactive and Compositional Systems
A different trend is to build systems that can combine multiple tools (i.e., existing vision-only or vision-and-language models), usually through ChatGPT or another LLM. In particular, these approaches aim to let the user interact with the LLM which is in charge of selecting the useful tools to carry out complex tasks. In this context, some solutions study how to prompt ChatGPT Wu et al. (2023a); Yang et al. (2023c) to invoke visual foundation models. GPT4Tools Yang et al. (2023a), instead, employs open-source LLMs such as LLaMA and OPT, that are fine-tuned with PEFT techniques to use tools for performing a wide range of visual tasks. Differently, Liu et al. (2023l) introduce more sophisticated user-chatbot interactions, through the incorporation of mouse-based pointing instructions on images or videos.
@@ Appendix E Interactive and Compositional Systems
VideoChat Li et al. (2023g) ${}^{\bigstar}$ EVA ViT-g Visual Dialogue, VQA, Captioning
Video-ChatGPT Maaz et al. (2023) ${}^{\bigstar}$ CLIP ViT-L Visual Dialogue, VQA, Captioning
Video-LLaMA Zhang et al. (2023b) ${}^{\bigstar}$ EVA ViT-g Visual Dialogue, Captioning, VQA, Audio Understanding
BT-Adapter Liu et al. (2023g) ${}^{\bigstar}$ CLIP ViT-L Visual Dialogue, Captioning, VQA, Retrieval
LLaMA-VID Li et al. (2023i) ${}^{\blacklozenge}$ EVA ViT-g Visual Dialogue, VQA, Captioning
PG-Video-LLaVA Munasinghe et al. (2023) ${}^{\bigstar}$ CLIP ViT-L Visual Dialogue, Captioning, VQA, Grounding
TimeChat Ren et al. (2023a) ${}^{\blacktriangle}$ EVA ViT-g Visual Dialogue, Captioning, Temporal Grounding, Highlight Detection
Vista-LLaMA Ma et al. (2023a) ${}^{\bigstar}$ EVA ViT-g Visual Dialogue, VQA, Captioning
@@
   mPLUG-DocOwl Ye et al. (2023a) ${}^{\blacktriangle}$ CLIP ViT-L Visual Dialogue, Captioning, VQA
   Kosmos-2.5 Lv et al. (2023) ${}^{\blacklozenge}$ Pix2Struct ViT-L Text Recognition, Image-to-Markdown Generation
   UReader Ye et al. (2023b) ${}^{\blacktriangle}$ CLIP VIT-L Visual Dialogue, VQA, Captioning, Information Extraction
   mPLUG-PaperOwl Hu et al. (2023) ${}^{\blacktriangle}$ CLIP ViT-L Visual Dialogue, VQA, Captioning, Diagram Analysis
   LLaMA-SciTune Horawalavithana et al. (2023) ${}^{\blacklozenge}$ CLIP ViT-L Visual Dialogue, VQA, Captioning, Diagram Analysis
   DocPedia Feng et al. (2023a) ${}^{\blacklozenge}$ Swin-B Visual Dialogue, VQA, Information Extraction
@@
   EmbodiedGPT Mu et al. (2023) ${}^{\bigstar}$ EVA ViT/g, RN50 Visual Dialogue, VQA, Captioning, Task Planning
   PaLM-E Driess et al. (2023) ${}^{\blacklozenge}$ ViT-22B Visual Dialogue, VQA, Captioning, Task Planning, Manipulation
@@
   PMC-VQA Zhang et al. (2023g) ${}^{\bigstar}$ PMC-CLIP RN50 VQA
   LLaVA-Med Li et al. (2023c) ${}^{\blacklozenge}$ CLIP ViT-L Visual Dialogue, VQA
   Qilin-Med-VL Liu et al. (2023f) ${}^{\blacklozenge}$ CLIP ViT-L Visual Dialogue, VQA
@@
   Dolphins Ma et al. (2023b) ${}^{\blacktriangle}$ CLIP ViT-L Visual Dialogue, VQA, Captioning, Traffic Condition Understanding
   DriveGPT4 Xu et al. (2023b) ${}^{\blacklozenge}$ CLIP ViT-L Visual Dialogue, VQA, Captioning
@@
   FoodLLM Yin et al. (2023b) ${}^{\blacktriangle}$ CLIP ViT-L Visual Dialogue, VQA, Nutrition Estimation, RES

### v2

@@ 3.2 Image Generation and Editing
While initial MLLMs excelled in extracting information from visual data, recent research included the generation of visual outputs. This advancement is realized through integrating MLLMs with image generation mechanisms, predominantly embodied by the Stable Diffusion (SD) Rombach et al. (2022) models. These models feature a denoising U-Net Ronneberger et al. (2015) architecture conditioned on textual or visual embeddings, through cross-attention layers. A complete list of the analyzed models is presented in Table 3.
@@
A different approach has been studied by Li et al. (2023f) which proposes to fine-tune the LLM by adding two special tokens (i.e., <start> and <end>), and directly encode the generated text between these two tokens using the text encoder in the SD model. Similarly, in Xia et al. (2023a) the LLM is trained to output detailed language-based generation prompts which are employed for generation or editing tasks. The U-Net is fine-tuned with longer and more detailed textual captions. Furthermore, in DreamLLM Dong et al. (2023) an alignment loss is eschewed in favor of a score distillation loss while keeping the U-Net frozen. Additional research endeavors have been conducted to introduce MLLMs in the field of image editing Fu et al. (2024); Huang et al. (2023c); Tang et al. (2023).
@@ 3.3 Other Modalities and Applications
Video Understanding. Although much of the research focuses on images, some works propose MLLMs specifically designed to handle video sequences. These models process video frames independently, using CLIP-based backbones to extract frame-level features which are then combined using pooling mechanisms Li et al. (2023j); Maaz et al. (2023) or Q-Former based solutions Li et al. (2023h); Ren et al. (2023a). The connection between visual features and the language model mainly follows the same trend as image-based MLLMs, with linear projections being the most common choice. However, there are also some attempts to develop video-specific adapters Liu et al. (2023g); Ma et al. (2023a) that can capture fine-grained temporal information. In addition to encoding video frames, some works Munasinghe et al. (2023); Zhang et al. (2023b) also employ audio features to enrich the representation of input video sequences. Furthermore, effective strategies for visual instruction tuning are also designed in the video domain Song et al. (2024), enabling more effective understanding of long video sequences.
Any-Modality Models. Almost all models described so far treat a single modality as input to the LLM. However, a significant body of work focuses on designing effective solutions that can handle multiple modalities. This is usually achieved by aligning multimodal features through Transformer blocks such as Q-Former Chen et al. (2023c); Panagopoulou et al. (2023) and Perceiver Zhao et al. (2023d), or by utilizing ImageBind to effectively extract features that are inherently multimodal Su et al. (2023). Images, videos, and audio are the most commonly treated modalities. Additionally, some works also effectively encode 3D data Yin et al. (2023d) and IMU sensor signals Moon et al. (2023). While all these solutions can manage multimodal inputs, approaches like NExT-GPT Wu et al. (2023c) and Unified-IO 2 Lu et al. (2023a) are also capable of generating outputs of different modalities.
Domain-Specific MLLMs. In addition to dealing with generic visual inputs, some research efforts are dedicated to developing MLLMs for specific domains and applications, either training the model starting from a pre-trained LLM or fine-tuning an existing MLLM with domain-specific data. Some examples are MLLMs designed for document analysis and text-intensive visual inputs Lv et al. (2023); Ye et al. (2023a), those proposed for embodied AI and robotics Driess et al. (2023); Mu et al. (2023), and those tailored for specific domains such as medicine Li et al. (2023d) and autonomous driving Xu et al. (2023c). A complete list of domain-specific MLLMs is reported in the supplementary.
@@ 4 Conclusion and Future Directions
In this survey, we have provided a comprehensive overview of the recent evolution of MLLMs, first focusing on how to equip LLMs with multimodal capabilities and then exploring the main tasks addressed by these models. Based on the analysis presented, in the following, we outline important open challenges and promising future research directions to further empower MLLMs.
Multimodal Retrieval-Augmented Generation. While retrieval-augmented generation (RAG) is a consolidated technique in LLMs Lewis et al. (2020); Asai et al. (2023), its application in MLLMs is still under-explored. We believe that the emergence of VQA datasets that require external retrieved knowledge Chen et al. (2023k); Mensink et al. (2023) may enable the development of MLLMs with RAG capabilities Hu et al. (2023b); Caffagni et al. (2024).
Correction of Hallucinations. Several studies Liu et al. (2023b); Zhu et al. (2023a) show that MLLMs tend to exhibit high hallucination rates, especially when generating longer captions. While some solutions are emerging to mitigate this problem Liu et al. (2023b); Wang et al. (2023a); Wu et al. (2023d); Yin et al. (2023b); Jing et al. (2023), understanding and correcting the underlying causes of hallucinations remains an important open challenge that is worth addressing to allow the application of these models in more critical contexts (e.g., medicine) and guarantee their accuracy and trustworthiness.
Prevent Harmful and Biased Generation. Ensuring the safety and fairness of large-scale models is of fundamental interest in the community. Recent works show that models trained on web-crawled data are prone to generate inappropriate and biased content. Although recent efforts are being made to reduce this phenomenon in text-to-image generative models Schramowski et al. (2023); Friedrich et al. (2023); Poppi et al. (2024), further exploration is needed to prevent the same behavior in MLLMs Pi et al. (2024).
@@
Limitations
This survey provides a comprehensive review of visual-based MLLMs. Although we have made a significant effort to include all relevant works available to the date of submission, the review might have missed some minor works, and might not have a complete coverage of MLLMs treating modalities that are different from the visual one. Additionally, given the space constraints required by the submission venue, we have restricted our explanations of existing approaches so as to include only the most relevant novelty points. We encourage the reader to refer to the original papers for further technical details and implementation notes.
Acknowledgments
@@
Appendix A Handling Images of Different Resolutions and Aspect Ratios
Most existing MLLMs perceive images in a low resolution and a fixed squared aspect ratio. Some works Liu et al. (2023d); You et al. (2023); Chen et al. (2023j) have demonstrated that adopting visual backbones trained on higher resolutions leads to fewer hallucinations and improved multimodal understanding abilities, translating into better performance over tasks that require fine-grained details. However, scaling an MLLM to arbitrary input resolutions and aspect ratios raises two important concerns: (i) the adaptation issue of switching from small images seen during training to larger ones at inference time and (ii) computational costs provided by the increased number of tokens in both the visual encoder and the LLM, given by the quadratic complexity of the attention-based architectures. In the following, we distinguish three different approaches to address these problems.
Positional-Encoding Interpolation. These models interpolate the positional encoding of their visual backbones, trained at low resolutions, to handle high-resolution images. While being simple, these methods are prone to adaptation issues. As a consequence, they partially mitigate this issue by performing at least one high-resolution training stage. To reduce the input sequence length to the LLM, and thus the computational cost, MiniGPT-v2 Chen et al. (2023e) and VILA Lin et al. (2023a) propose to project multiple visual tokens together into the same token within the embedding space of the LLM. For the same reason, mPLUG-Owl2 Ye et al. (2023d) and Qwen-VL Bai et al. (2023b) compress the visual features into fixed-length sequences, independent of the resolution, using learnable queries. The latter further saves computation in most layers of the ViT backbone due to a window attention mechanism.
Sub-Images Slicing. To avoid the adaptation issue, some methods propose to slice a high-resolution image into multiple sub-images of fixed size according to the native resolution of their visual encoder. Then, each sub-image is processed independently by the visual backbone, along with the whole image downsized at the same resolution, and the features are concatenated to obtain the global representation. SPHINX Lin et al. (2023b) divides the input image in a squared grid of sub-images (i.e., $2\times 2$ or $3\times 3$) at the training resolution of the visual backbone. Moreover, to handle rectangular aspect ratios, SPHINX pads the image to reach the desired square size. For extreme aspect ratios, the padding leads to sub-images which are only composed of padding. Hence, in SPHINX-X Gao et al. (2024) a skip token is introduced to replace noisy tokens associated with only padding sub-images and reduce the sequence length provided to the LLM, increasing efficiency. Similarly, LLaVA-NeXT Liu et al. (2024b) ignores sub-images composed only of padding and handles different shapes of grids, by introducing a special token that indicates when a row of sub-images ends. Monkey Li et al. (2023l) uses a Perceiver-like resampler to extract fixed-length sequences from each sub-image and trains on an image-text dataset curated by several vision expert models integrated by ChatGPT. InfiMM-HD Liu et al. (2024a) proposes a dynamic resolution adaptation training stage to increase the image size up to 1,344 pixels. It employs gated cross-attention layers (as in Flamingo) to inject the visual features into the LLM, without increasing the input sequence length. LLaVA-UHD Xu et al. (2024) finds the optimal partitioning scheme leading to sub-images that most resemble the native resolution and aspect ratio of the visual encoder. The number of visual tokens is compressed through a Perceiver-like adapter and employs a spatial schema in such a way that the LLM can understand the grid of sub-images.
Others. Another solution, namely OtterHD Li et al. (2023a), can seamlessly deal with any resolution or aspect ratio, as it directly feeds large image patches of $30\times 30$ pixels to the LLM, without the need for a visual encoder. LLaVA-HR Luo et al. (2024), instead, introduces a mixture-of-resolution adaptation to fuse into the ViT layers high-resolution features extracted with a CNN and low-resolution ones produced by the ViT itself.
Appendix B Additional Training Data
@@ Appendix C Evaluation
C.1 Standard Benchmarks
One of the most important skills of MLLMs is their ability to effectively answer questions based on the given input image. This ability is quantitatively evaluated across several visual question-answering datasets, measuring the accuracy (Antol et al., 2015) of the answers provided by the MLLM.
@@ C.1 Standard Benchmarks
C.2 MLLM-Specific Benchmarks
@@
POPE Li et al. (2023k) is a valuable benchmark for evaluating object hallucination challenges within MLLMs. This dataset encompasses several distinct subsets, namely random, popular, and adversarial, which are generated utilizing a variety of sampling methodologies. Cumulatively, it is a binary classification query dataset that comprises 8,910 entries, facilitating comprehensive investigations into the phenomenon of object hallucination within the context of MLLMs.
@@
SEED-Bench (SEED) Li et al. (2023c) is specifically designed to evaluate LLMs and MLLMs across 12 dimensions spanning from scene understanding to OCR and action recognition. The benchmark consists of 19k multiple-choice questions written by human annotators.
@@
C.3 Visual Grounding Evaluation
@@
Visual Genome Krishna et al. (2017) connects structured image concepts to language and comprises 108,077 images along with detailed descriptions of all objects present in them, providing 5.4M region descriptions and 1.7M visual question-answer pairs. This dataset is typically used for region-level captioning and question-answering.
Visual7W Zhu et al. (2016) is a visual question-answering dataset that combines textual descriptions with image regions through object-level grounding. It comprises 328k question-answer pairs on 47k COCO images, together with 1.3M human-generated multiple-choice and more than 560k object groundings from 36,579 categories.
@@
C.4 Image Generation and Editing Evaluation
@@
PaLI Chen et al. (2023j) TPUv4 1,024
@@
Appendix D Computational Requirements
@@
VideoChat Li et al. (2023h) StableVicuna-13B EVA ViT-g Visual Dialogue, VQA, Captioning
Video-ChatGPT Maaz et al. (2023) Vicuna-7B CLIP ViT-L Visual Dialogue, VQA, Captioning
Video-LLaMA Zhang et al. (2023b) Vicuna-7B EVA ViT-g Visual Dialogue, Captioning, VQA, Audio Understanding
BT-Adapter Liu et al. (2023g) Vicuna-7B CLIP ViT-L Visual Dialogue, Captioning, VQA, Retrieval
LLaMA-VID Li et al. (2023j) Vicuna-13B EVA ViT-g Visual Dialogue, VQA, Captioning
PG-Video-LLaVA Munasinghe et al. (2023) LLaVA-1.5-13B CLIP ViT-L Visual Dialogue, Captioning, VQA, Grounding
TimeChat Ren et al. (2023a) LLaMA-2-7B EVA ViT-g Visual Dialogue, Captioning, Temporal Grounding, Highlight Detection
Vista-LLaMA Ma et al. (2023a) Vicuna-7B EVA ViT-g Visual Dialogue, VQA, Captioning
@@
Appendix E Additional Details on Other Modalities and Applications
Video Understanding. As a complement of Sec. 3.3, we report in Table 12 a summary of the main characteristics of video-based MLLMs. For each model, we indicate the LLM used as starting point, which in some cases is initialized with the parameters of a pre-trained MLLM, the visual encoder, and the main tasks and capabilities of the MLLM. Additionally, we specify whether the LLM is kept frozen, is entirely fine-tuned, or is fine-tuned with PEFT-based strategies.
3D Understanding. MLLMs are also applied to 3D data for solving complex tasks like 3D VQA, 3D conversation, and 3D dense captioning. Differently from standard visual encodings which exploit 2D pre-trained embeddings, in the context of 3D data, appropriate strategies are designed to project them to the LLM space. In 3D-LLM Hong et al. (2023), 3D scenes are rendered in different views and 3D features are built using an EVA-CLIP backbone connected to a fine-tuned BLIP-2 model. Similarly, Xu et al. (2023b) employ a pre-trained Point-BERT Yu et al. (2022) as 3D encoder and conducts a two-stage training that initially aligns the input features via an MLP projection layer, and then performs an instruction tuning phase of the model. Differently, in Point-Bind Guo et al. (2023), 3D point-clouds are aligned with ImageBind Girdhar et al. (2023) and by leveraging I2P-MAE Zhang et al. (2023f) as 3D encoder. This alignment allows the introduction of new tasks such as any-to-3D generation and 3D embedding-space arithmetic. Recently, LL3DA Chen et al. (2023g) introduces the Interactor3D module, which consists of a frozen 3D scene encoder, a visual prompt encoder, and a Q-Former to address 3D captioning and VQA.
@@ Appendix E Additional Details on Other Modalities and Applications
mPLUG-DocOwl Ye et al. (2023a) mPLUG-Owl-7B CLIP ViT-L Visual Dialogue, Captioning, VQA
Kosmos-2.5 Lv et al. (2023) Magneto-1.3B Pix2Struct ViT-L Text Recognition, Image-to-Markdown Generation
UReader Ye et al. (2023b) mPLUG-Owl-7B CLIP VIT-L Visual Dialogue, VQA, Captioning, Information Extraction
mPLUG-PaperOwl Hu et al. (2023a) mPLUG-Owl-7B CLIP ViT-L Visual Dialogue, VQA, Captioning, Diagram Analysis
LLaMA-SciTune Horawalavithana et al. (2023) LLaMA-13B CLIP ViT-L Visual Dialogue, VQA, Captioning, Diagram Analysis
DocPedia Feng et al. (2023) Vicuna-7B Swin-B Visual Dialogue, VQA, Information Extraction
@@
EmbodiedGPT Mu et al. (2023) LLaMA-7B EVA ViT/g, RN50 Visual Dialogue, VQA, Captioning, Task Planning
PaLM-E Driess et al. (2023) PaLM-540B ViT-22B Visual Dialogue, VQA, Captioning, Task Planning, Manipulation
@@
PMC-VQA Zhang et al. (2023h) PMC-LLaMA-7B PMC-CLIP RN50 VQA
LLaVA-Med Li et al. (2023d) LLaVA-7B CLIP ViT-L Visual Dialogue, VQA
Qilin-Med-VL Liu et al. (2023f) CN-LLaMA2-13B CLIP ViT-L Visual Dialogue, VQA
@@
Dolphins Ma et al. (2023b) OpenFlamingo-7B CLIP ViT-L Visual Dialogue, VQA, Captioning, Traffic Condition Understanding
DriveGPT4 Xu et al. (2023c) LLaMA-2-7B CLIP ViT-L Visual Dialogue, VQA, Captioning
@@
FoodLLM Yin et al. (2023c) LISA-7B CLIP ViT-L Visual Dialogue, VQA, Nutrition Estimation, RES
@@
Any-Modality Models. Several studies focus on extending the reasoning capabilities of the MLLMs by including multiple modalities, such as video, 3D, and audio. A line of research investigates the usage of dedicated pathways to input the different modalities to the LLM. UniVAL Shukor et al. (2023) maps the features from each modality encoder into the shared representation space of the LLM through dedicated linear projections. X-LLM Chen et al. (2023c) leverages Q-Former interfaces for the image and video modalities, interpreting the video as a sequence of independent frames, each one encoded as an image. For the speech modality, it uses a C-Former interface that compresses the feature sequence from the speech encoder into token-level embeddings. X-InstructBLIP Panagopoulou et al. (2023) and ChatBridge Zhao et al. (2023d) propose to freeze both the modality encoders and the LLM and to leverage, respectively, dedicated Q-Former or Perceiver adapters for each modality. To maximize feature compatibility, AnyMAL Moon et al. (2023) uses an encoder that has already been aligned to a text embedding space for each modality, including also IMU signals, and a dedicated adapter, which is a Perceiver for the visual modality and linear layers for the others. On the other hand, PandaGPT Su et al. (2023) and NExT-GPT Wu et al. (2023c) exploit a single frozen multimodal encoder (i.e., ImageBind) to extract features from different modalities. OneLLM Han et al. (2024) builds a unified universal encoder and a universal projection module by mixing multiple image projection modules and a modality router to align input signals with language. CAT Ye et al. (2024) adds a clue aggregator to aggregate question-aware audio-visual hidden features and produce clue tokens that are provided to the LLM.
In addition to handling different modalities in input to the LLM, some works investigate the generation of outputs of different modalities. For example, NExT-GPT Wu et al. (2023c) introduces signal tokens in the LLM that indicate whether the diffusion-based decoder for a specific modality has to be activated. Moreover, the signal tokens are provided to a transformer-based output projector to condition the generation. Similarly, M2UGen Hussain et al. (2023) handles the music modality by using the LLM output corresponding to signal tokens, along with unimodal music features from a music encoder, to condition the generation of an audio encoder. LLMBind Zhu et al. (2024) indicates the conditioning text to generate image, video, or audio by wrapping it in special tokens. Thus, this text is provided to the corresponding modality-specific diffusion model. Unified-IO2 Lu et al. (2023a) uses VQ-GAN decoders for both image and audio modalities to decode output discrete tokens and can generate surface normals, depth, and segmentation masks for the input images. AnyGPT Zhan et al. (2024) interprets all the continuous non-text modalities as discrete tokens in both input and output, using, respectively, multimodal tokenizers and de-tokenizers. To enable the 3D modality, LAMM Yin et al. (2023d) introduces a novel instruction tuning dataset and benchmark that comprise both image-text and point cloud-text instruction-response pairs, covering a wide range of 2D and 3D tasks.
Interactive and Compositional Systems. A different trend is to build systems that can combine multiple tools (i.e., existing vision-only or vision-and-language models), usually through ChatGPT or another LLM. In particular, these approaches aim to let the user interact with the LLM which is in charge of selecting the useful tools to carry out complex tasks. In this context, some solutions study how to prompt ChatGPT Wu et al. (2023a); Yang et al. (2023c) to invoke visual foundation models. GPT4Tools Yang et al. (2023a), instead, employs open-source LLMs such as LLaMA and OPT, that are fine-tuned with PEFT techniques to use tools for performing a wide range of visual tasks. Differently, Liu et al. (2023l) introduce more sophisticated user-chatbot interactions, through the incorporation of mouse-based pointing instructions on images or videos.
Domain-Specific MLLMs. Finally, in Table 13 we summarize the main characteristics of domain-specific MLLMs, also in this case indicating for each model the LLM used as starting point.